OpenStudy (anonymous):

Which of the following inequalities matches the graph of a solid line through the points (0,1) and (2, -3) with shading above the line? a. y ≥ -2x + 1 b. y > -2x + 1 c. y ≤ -2x + 1

OpenStudy (jdoe0001):

so what do you think is the slope of the line that passes through -> (0,1) and (2, -3) ?

OpenStudy (anonymous):

slope is -2/1?

OpenStudy (jdoe0001):

hmm, wait a second, I notice all right sides are the same, nevermind the slope then

OpenStudy (jdoe0001):

yes is -2, but all have -2 so well, anyhow a SOLID LINE in inequality graphing means "y is GREATER THAN" or "y is LESS THAN"

OpenStudy (jdoe0001):

as opposed to GREATER THAN OR EQUALS TO, which means a "dashed line"

OpenStudy (anonymous):

oh I missed choice d. The correct inequality is not listed. That's what I think the answer is

OpenStudy (jdoe0001):

so, what's D?

OpenStudy (jdoe0001):

because since it's a SOLID line, you can pretty much rule out A and B

OpenStudy (jdoe0001):

I mean A and C

OpenStudy (anonymous):

d. The correct inequality is not listed.

OpenStudy (jdoe0001):

ohh that's D: not listed, darn didn't read correctly heheh

OpenStudy (jdoe0001):

well, ruling out A and C only leaves B since all A B and C are using the same form for the right-side, then I'd gather that's the equation, and so the choice will be the obvious one

OpenStudy (anonymous):

choice d?

OpenStudy (jdoe0001):

h..... wait... darn I got mixed

OpenStudy (jdoe0001):

hold the mayo, a solid line means \(\bf \ge\) or \(\bf \le\)

OpenStudy (jdoe0001):

so is either A or C, man dohhh

OpenStudy (anonymous):

Ack! I hate these things!

OpenStudy (jdoe0001):

my bad, that was my mix up though

OpenStudy (jdoe0001):

lemme get the line

OpenStudy (jdoe0001):



OpenStudy (jdoe0001):

there, as you can see, is a declining function, a negative slope -2 so now let's test around some points NOT IN THE LINE say let's test for say (-2, 1) is not the line let's plug those values in A) 1 ≥ -2(-2) + 1 1 ≥ 5 so we know that \(\bf 1 \cancel{ \ge } 5\) so it's not A

OpenStudy (jdoe0001):

now let's test C, our other candidate same point (-2, 1) not in the line 1 ≤ -2(-2) + 1 1 ≤ 5 so we do know that indeed \(\large \bf 1 \le 5\)

OpenStudy (jdoe0001):

so, that's your choice

OpenStudy (jdoe0001):

sorry on the mix up solid lines mean GREATER/LESS OR EQUAL dashed lines mean GREATER/LESS

OpenStudy (anonymous):

wow thanks!!

OpenStudy (jdoe0001):

yw